OpenStudy (anonymous):

A cone has a volume of 12 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of? a. 6 in3 b. 24 in3 c. 36 in3 d. 48 in3

OpenStudy (anonymous):

heyyyy its @SolomonZelman

OpenStudy (ranga):

Volume of a cone = 1/3(pi)(r^2)h Volume of a cylinder with the same dimension is pi(r^2)h Take the ratio of the two volumes: cylinder volume / cone volume first. What do you get?

OpenStudy (anonymous):

how do i get the height

OpenStudy (anonymous):

would it be 36? @ranga

OpenStudy (ranga):

Yes, 36. When you take the ratio the height will cancel out. cylinder volume / cone volume = pi(r^2)h / 1/3(pi)(r^2)h = 3 So the cylinder will have three times the volume of the cone. So 3 x 12 = 36 cubic inches.